Larry is three times as old as Moe, and Curly is sixteen years younger than Larry, Larry's age was twice the sum of the ages of Moe and Curly. Find each man's age.
Answer by fractalier(2101)

Well, this one works out kind of funny, but it works out. First, write the equations from the info given. Thus we have
L = 3M
C = L - 16
L = 2(M + C)
Now substitute into that third equation values for M and C to get
L = 2[(L/3) + (L - 16)]
L = (8/3)L - 32 and
(-5/3)L = -32 so that
L = 96/5 = 19.2 yrs From that
C = 16/5 = 3.2 yrs and
M = 32/5 = 6.4 yrs